You put half of your money in a stock portfolio that has an expected return of 14% and a standard deviation of 24%. You put the rest of your money in a risky bond portfolio that has an expected return of 6% and a standard deviation of 12%. The stock and bond portfolios have a correlation of .55 and TB or Rf is 2%. What is the SR of the resulting portfolio?
Answer #1

Expected Return = 0.50(0.14) + 0.50(0.06) = 10.00%

Standard Deviation = [(0.50)2(0.24)2 + (0.50)2(0.12)2 + 2(0.50)(0.50)(0.24)(0.12)(0.55)]1/2

Standard Deviation = 16.10%

Sharpe Ratio = (Rp - Rf)/Standard Deviation

Sharpe Ratio = (0.10 - 0.02)/0.1610

Sharpe Ratio = 0.50